OpenStudy (anonymous):

# How many possible positive, negative and complex solutions are there in this equation x^3-10x^2+21x-210

OpenStudy (anonymous):

are they asking you to find the exact number of solutions that are positive negative and complex...or do they want to know how many solutions there can be?

OpenStudy (anonymous):

Equals zero? Total number of real and complex roots is three, based on the degree of the function. At least one will be real.

OpenStudy (anonymous):

Factor by grouping to find out the particulars of this problem.

sam (.sam.):

3

OpenStudy (amistre64):

its a Dsign question

OpenStudy (anonymous):

Yeah factor and find as many solutions...if there are still some solutions left that you could not find through factoring then the leftover solutions will be complex

sam (.sam.):

x^(3)-10x^(2)+21x-210=0 (x^(2)+21)(x-10)=0 x^(2)+21=0 x-10=0 x=i sqrt(21),-i sqrt(21),10

OpenStudy (amistre64):

x^3 -10x^2 +21x -210 1 | 1 | 1 | 0 ; 3 sign changes in the normal manner there are possibly 3 or 1 positive zeros

OpenStudy (amistre64):

swap signs on odd powers -x^3 -10x^2 -21x -210 0 | there are no sign changes in the negative version; there are no negative roots

OpenStudy (amistre64):

there are possibly 2 complex roots which is why we had to step the first one by 2

OpenStudy (anonymous):

\[x^3-10x^2+21x-210=0\rightarrow x^2(x-10)+21(x-10)\]\[=(x^2+21)(x-10)=0\rightarrow x \in \left\{ 10,\pm i \sqrt{21} \right\}\]

OpenStudy (anonymous):

okay so there are possibly 3 or 1 positive zeros and 2 complex but no negative?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

youre welcome